OpenStudy (anonymous):

Identify the vertex and the y-intercept of the given parabola: y=-3(x+2)^2+5

jhonyy9 (jhonyy9):

y=-3(x^2 +2x +4) +5 y=-3x^2 -6x -12 +5 y= -3x^2 -6x -7

OpenStudy (saifoo.khan):

@jhonyy9 , they asked for vertex. not the solutions/

OpenStudy (anonymous):

I figured out the vertex, I just don't know the y-solution. It's either -7 or 7. I'm not sure though.

OpenStudy (saifoo.khan):

y=-3(x+2)^2+5 In this problem, The vertex is (-2, 5) We found this as we have to change the sign for the term with "x" and we write the outside co-efficent as it is. For y-intercpet, make x = 0 -3(0 + 2)^2 + 5 -3(2)^2 + 5 -3(4) + 5 -12 + 5 -7

OpenStudy (anonymous):

Ah thank you so much!!!!! :)

OpenStudy (saifoo.khan):

No problem! :)